# Evaluating infectious disease forecasts with allocation scoring rules

### 24 Feb 2024 v2

@@ 2 The Allocation Score
We begin with an informal description of the allocation score and some examples illustrating its key characteristics in section 2.1. In section 2.2 we develop the allocation score more carefully, using a decision theoretic procedure for deriving proper scoring rules. (See appendix B for a definition of a proper scoring rule and a more technical discussion of the procedure). In section 2.3, we note that another group of common scores including the quantile score, WIS, and CRPS, can also be derived from decision theoretic foundations —starting from a different decision making context.
@@ 2.1 Overview of Allocation Scoring
Suppose we have a forecast $F$ for need in two locations with $F_{1}=\mathrm{Exp}(1/\sigma_{1})$ and $F_{2}=\mathrm{Exp}(1/\sigma_{2})$, where $\sigma_{1}=1$ and $\sigma_{2}=4$. The means of these distributions are given by the scale parameters $\sigma_{i}$. When the marginal forecasts are exponential distributions, it can be shown that the optimal allocation divides the available resources among the locations proportionally to the scale parameters $\sigma_{i}$ (see appendices D). If $K=5$ units of the resource are available, the optimal allocation according to $F$ would be 1 and 4 resource units in locations 1 and 2, respectively. Increasing $K$ to $10$ changes this optimal allocation to 2 and 8. Figure 1 illustrates the situation.
@@ 2.2.1 The decision theoretic setup for forecast evaluation
This is a general procedure that may be applied in settings where it is possible to specify a quantitative loss function. We call a scoring rule obtained from this procedure a Bayes scoring rule (noting that to our knowledge, this terminology is not standard). In appendix B we review the result from the scoring rule literature (e.g., Dawid (2007); Gneiting and Raftery (2007)) that Bayes scoring rules are proper by construction.
@@ 2.2.2 The allocation score for a fixed resource constraint
In our setting an action is a vector $x=(x_{1},\ldots,x_{N})$ specifying the amount of resources allocated to each of $N$ locations. We require that $0\leq x$, i.e., that each $x_{i}$ is non-negative, and that the total allocation across all locations equals the amount of available resources, $K$: $\sum_{i=1}^{N}x_{i}=K$. The set $\mathcal{X}$ consists of all possible allocations that satisfy these constraints. The eventually realized resource need in each location is denoted by $y=(y_{1},\ldots,y_{N})$; this is a realization (unknown at the time of decision making) of the random vector $Y=(Y_{1},\ldots,Y_{N})$. A forecast of the random need $Y$ for each location is written as $F=(F_{1},\ldots,F_{N})$. We assume need is non-negative and the forecasts reflect that, i.e. the support of each $F_{i}$ is a subset of $\mathbb{R}^{+}$. Finally, we assume that each unmet unit of need in any location incurs a fixed loss of $L>0$, so that if the selected resource level $x_{i}$ in location $i$ is less than the realized need $y_{i}$, a loss of $L\cdot(y_{i}-x_{i})$ results. A variety of extensions to this setup are possible; for example, we might account for storage costs for resources that go unused, allow for a different loss per unit of unmet need in each location, or account for resource transportation costs. In this work, we choose to keep the loss function relatively simple to focus on the core ideas.
@@
In appendix D we show that the components of the Bayes act are quantiles $x_{i}^{F,K}=F_{i}^{-1}(\tau^{F,K})$ at a probability level $\tau^{F,K}$ that depends on the forecast $F$ and the resource constraint $K$, but is shared across all locations. This probability level is the level at which the resource constraint is satisfied: $\sum_{i=1}^{N}F_{i}^{-1}(\tau^{F,K})=K$. This tells us that in order to allocate optimally (according to $F$), we must divide resources among the locations so that there is an equal forecasted probability in every location that the allocation is sufficient to meet resource need. This solution to the allocation problem is well-known in inventory management and is often attributed to Hadley and Whitin (1963).
@@
We can now use the Bayes act to define a proper scoring rule for the probabilistic forecast $F$. Consider first the “raw” score defined as
@@ 2.3 Connections to Other Scores
The weighted interval score (WIS) was proposed in 2020 as a way to score forecasts that were being made in the early stages of the COVID-19 pandemic (Bracher et al., 2021); equivalent scores had also been used in previous forecast evaluation efforts (e.g., Hong et al., 2016). The WIS is a proper scoring rule for forecasts that use a set of quantiles to represent a probabilistic forecast distribution. While pointing a reader interested in more mathematical detail to Bracher et al. (2021), we note simply that the WIS is a weighted sum of interval scores at different probability levels (e.g., 50% prediction intervals, 80% PIs, 95% PIs, etc…). Larger interval scores indicate less skillful forecasts. An interval score consists of (a) the width of the interval, with larger intervals receiving higher scores (higher scores indicate less accuracy), and (b) a penalty if the interval does not cover the eventual observation, which increases the further away the interval is from the observed value. Equivalently, the WIS can also be characterized as a weighted sum of quantile scores for each individual predictive quantile. The quantile score for a particular quantile level assigns an asymmetric penalty to predictions that are too high or too low, with the relative sizes of the penalties set so that in expectation the score is minimized by the given quantile of the distribution. The most commonly used version of WIS is one that uses an equal weighting of all quantile levels, in which case WIS approximates the continuous ranked probability score (CRPS), a commonly used score for probabilistic forecasts. It is important to note that this weighting was proposed because the resulting score approximates the CRPS, and not because it aligned with any particular public health decision-making rationale.
@@ 3.2 Evaluation metrics
As described above, predictions were submitted to the Forecast Hub in the form of a set of 23 quantiles of the predictive distribution. The WIS can be directly calculated from these quantiles. However, calculation of the AS requires that the full cumulative distribution functions of the forecast distributions for each location are available. For the purpose of this analysis, to calculate the allocation score we approximated the full cumulative distribution functions based on the provided quantiles. On the interior of the provided quantiles we used a monotonic cubic spline to interpolate the quantiles, and in the lower and upper tails we used normal distributions with parameters selected so as to match the two lowest and two highest quantiles (see appendix F for further details). In appendix G, we show that if this evaluation procedure had been specified prospectively, the resulting score would be proper, —but that a post hoc application of this procedure is improper. We use the procedure here to illustrate the properties of the AS, and note that a collaborative forecasting hub interested in using the AS for evaluation could circumvent propriety issues by communicating the definition of the AS to forecasters and then collecting allocations at specified resource levels as part of forecast submissions.
@@ 3.5.3 Metrics were not consistently correlated over time
We can gain some insight into the discordance between MAS and taMWIS by examining the correlation between time-specific AS and MWIS values of the various models in Figure 5. We find, for example, no clear association between the consistent and high MWIS rank of the JHUAPL-SLPHospEns model and its highly variable AS rank. This contrasts with a clearly positive correlation between MWIS and AS ranks in other models such as Karlen-pypm and USC-SI_kJalpha. We also can see in more detail the consistently high performance COVIDhub-ensemble on both scores. And while it did not submit forecasts for enough weeks to be included in Table 2, CMU-TimeSeries is an interesting example of a model that performed consistently well on AS but had only middling MWIS ranks.
@@ 3.5.4 Integrated allocation score across values of K
The integrated allocation score (IAS) summarizes allocation scores (AS) across a range of possible values of the constraint ($K$), possibly assigning higher weights to more likely values of $K$ (Section 2.2.3). IAS was computed for two weight distributions on $K$, one uniform across the entire range and the other a symmetric distribution around 15,000 (the orange and blue shaded areas in Figure 6A, respectively). Both versions of the IAS were correlated with the original AS for $K=15,000$, with the higher correlation coming from the weighting centered at $K=15,000$ (Figure 6B). Model rankings based on the AS and the centered IAS were roughly similar, with the top and bottom three approaches being the same for both scores (Table 1).
@@ 4 Discussion
Probabilistic forecasts of infectious disease outbreaks have typically been evaluated using well-known proper scoring rules such as the LS, CRPS, and WIS. Often models are ranked according to such a score, but without reference to any underlying decision problem from which the score might be derivable as a Bayes scoring rule or be otherwise motivated. We argue that the value of collaborative forecasting efforts, such as the COVID-19 Forecasting Hub, could be enhanced by including evaluations of forecasts with scoring rules that directly consider the success of forecasts in supporting specific public health decisions. As evidence for this possibility, we have demonstrated how tying forecast evaluation to a specific decision problem (via the allocation score) can yield model rankings that differ substantively from those based a current standard measure of forecast accuracy, the WIS, which does not refer to any decision problem that has been clearly connected to outbreak response. This result aligns with findings in other fields, especially those where the monetary consequences of using a forecast can be clearly juxtaposed with standard accuracy measures of the forecast (Leitch and Tanner, 1991; Murphy, 1993; Cenesizoglu and Timmermann, 2012). Additionally, we show that an existing ensemble forecast approach was the only method to outperform (in terms of this new evaluation method) two benchmark allocation approaches in a hypothetical application, suggesting that there is room for innovation of current epidemiological modeling techniques that might be spurred on by a reorientation of the field toward practical decision problems. For example, in order to surpass benchmark allocation scores, forecasters need to devote effort to capturing the relative magnitude of future resource need across different locations. Success at this task is not, however, directly rewarded by the WIS.
@@
In practice, epidemiological forecasts are often directed at a diverse set of end-users. It may be easy for some of these users to summarize the consequences of how they use a forecast with a numerical loss function, but for others, such as officials deciding how best to update public “situational awareness” of an outbreak, this may be essentially impossible. Even those forecast uses that are easily framed as expected loss minimization may differ enough that no single score would be appropriate for all users. Ideally, targeted forecasting tools could be developed through close collaboration between modelers and public health officials. However, this may only be possible in settings with sufficient staffing on both an analytics and a public health team. Increasingly, collaborative modeling hubs are being used to generate “one-size-fits-all” forecasts for many locations at once. But it could still be beneficial in these settings to expand the set of scoring methods used by the hub to include scores, such as the AS, which are based on specific public health decision problems. This could help end users better understand the value of available forecasts as inputs to their particular decision making contexts. (We note that issues of propriety are raised when forecasts are evaluated with scoring rules for which they were not elicited or via parameters imputed by a hub after submission, both of which are done in our application example. We consider these issues briefly in appendices F and G.)
@@
There are several important limitations to the current work. The allocation score we developed here does not directly account for important considerations such as fairness or equity of allocations, or more broadly, individual and group preference relations that are difficult or even impossible to encode into a loss or utility function. Ambiguity aversion (in the sense of Ellsberg (1961)), for example, seems especially relevant to the use of forecasts in outbreak response, where there can be immense social and political pressure on public health officials to maintain a transparent base of evidence for their policy choices and recommendations. The proposed framework also does not attempt to capture the broader context of decision making. For example, in practice it may be possible to increase the resource constraint $K$ by shifting funding from other disease mitigation measures. Finally, we emphasize that we opted at this stage of development to explicitly rule out situations where a “successful” epidemiological forecast could lead to policy decisions that change the distribution of the predicted outcome $Y$. Our framework would need considerable enhancements before being applicable to forecast evaluation beyond horizons for which causal feedback can be neglected.
@@
An opportunity for further investigation is to more carefully evaluate whether forecasts add value to “standard operating procedures” already in place for making public health decisions. In the context of allocations, such a procedure might extrapolate need from current observed need and population levels (similarly to the two benchmark approaches presented above), with adjustments based on other political or real-world considerations. For example, in many settings public health stakeholders will make decisions after synthesizing information from a variety of quantitative and qualitative sources coupled with expert judgment. The allocation score presented in this work does not directly measure whether a given forecast adds useful information to such an existing decision-making process. While the scoring procedures as presented do not directly address this question, they could be modified (say, by comparison to a ‘baseline’ model or expert-elicited allocations in the absence of forecast data) to quantify the benefit of using a forecast to inform a specific decision.
@@ Acknowledgements
Appendix A Introduction
We address some technical and methodological points from the main text. We begin in section B by defining proper scoring rules and showing that the allocation score is proper. In section C, we formalize the concept of a shortage of resources, give some key results about expected resource shortages under a distribution characterizing uncertainty about (future) levels of resource need, and illustrate how quantiles arise as the Bayes act in a decision making problem about the quantity of a resource to purchase. Section D gives a justification for the result that the Bayes act for the allocation problem is given by a vector of quantiles of the forecast distributions in each location at a shared probability level, which was stated in section 2.2.2. We describe the algorithm that we use to compute allocations given a forecast distribution in each location in section E. In section F, we describe the methods for approximating a full distribution from a set of quantiles, implemented in the R package distfromq, that we used to support computation of allocation scores from quantile forecasts that were submitted to the US COVID-19 Forecast Hub for the application in section 3. Finally, in section G we examine implications for propriety of an analysis that uses summaries of forecasts (such as predictive quantiles) rather than the full forecast distributions for computation of allocation scores.
Appendix B Proper scoring rules
@@
B.1 The allocation score is proper

### 04 Mar 2024 v3

@@ Abstract
Keywords: public health, forecast evaluation, proper scoring rules, resource allocation, epidemiology
@@ 2 The Allocation Score
We begin with an informal description of the allocation score and some examples illustrating its key characteristics in section 2.1. In section 2.2 we develop the allocation score more carefully, using a decision theoretic procedure for deriving proper scoring rules. (See appendix A for a definition of a proper scoring rule and a more technical discussion of the procedure). In section 2.3, we note that another group of common scores including the quantile score, WIS, and CRPS, can also be derived from decision theoretic foundations —starting from a different decision making context.
@@ 2.1 Overview of Allocation Scoring
Suppose we have a forecast $F$ for need in two locations such that $F_{1}$ and $F_{2}$ have exponential marginal distribution with scale parameters $\sigma_{1}=1$ and $\sigma_{2}=4$. The means of these distributions are given by the scale parameters $\sigma_{i}$. When the marginal forecasts are exponential distributions, it can be shown that the optimal allocation divides the available resources among the locations proportionally to the scale parameters $\sigma_{i}$ (see appendix B). If $K=5$ units of the resource are available, the optimal allocation according to $F$ would be 1 and 4 resource units in locations 1 and 2, respectively. Increasing $K$ to $10$ changes this optimal allocation to 2 and 8. Figure 1 illustrates the situation.
@@ 2.2.1 The decision theoretic setup for forecast evaluation
This is a general procedure that may be applied in settings where it is possible to specify a quantitative loss function. We call a scoring rule obtained from this procedure a Bayes scoring rule (noting that to our knowledge, this terminology is not standard). In appendix A we review the result from the scoring rule literature (e.g., Dawid (2007); Gneiting and Raftery (2007)) that Bayes scoring rules are proper by construction.
@@ 2.2.2 The allocation score for a fixed resource constraint
In our setting an action is a vector $x=(x_{1},\ldots,x_{N})$ specifying the amount of resources allocated to each of $N$ locations. We require that each $x_{i}$ is non-negative, and that the total allocation across all locations equals the amount of available resources, $K$: $\sum_{i=1}^{N}x_{i}=K$. The set $\mathcal{X}$ consists of all possible allocations that satisfy these constraints. The eventually realized resource need in each location is denoted by $y=(y_{1},\ldots,y_{N})$; this is a realization (unknown at the time of decision making) of the random vector $Y=(Y_{1},\ldots,Y_{N})$. A forecast of the random need $Y$ for each location is written as $F=(F_{1},\ldots,F_{N})$. We assume need is non-negative and that the forecasts reflect that, i.e. the support of each $F_{i}$ is a subset of $\mathbb{R}^{+}$. Finally, we assume that each unmet unit of need in any location incurs a fixed loss of $L>0$, so that if the selected resource level $x_{i}$ in location $i$ is less than the realized need $y_{i}$, a loss of $L\cdot(y_{i}-x_{i})$ results. A variety of extensions to this setup are possible; for example, we might account for storage costs for resources that go unused, allow for a different loss per unit of unmet need in each location, or account for resource transportation costs. In this work, we choose to keep the loss function relatively simple to focus on the core ideas.
@@
In appendix B we show that the components of the Bayes act are quantiles $x_{i}^{F,K}=F_{i}^{-1}(\tau^{F,K})$ at a probability level $\tau^{F,K}$ that depends on the forecast $F$ and the resource constraint $K$, but is shared across all locations. This probability level is the level at which the resource constraint is satisfied: $\sum_{i=1}^{N}F_{i}^{-1}(\tau^{F,K})=K$. This tells us that in order to allocate optimally (according to $F$), we must divide resources among the locations so that there is an equal forecasted probability in every location that the allocation is sufficient to meet resource need. This solution to the allocation problem is well-known in inventory management and is often attributed to Hadley and Whitin (1963).
@@
We can now use the Bayes act to define a proper scoring rule for the probabilistic forecast $F$. Consider first the raw score defined as
@@ 2.3 Connections to Other Scores
The weighted interval score (WIS) was proposed in 2020 as a way to score forecasts that were being made in the early stages of the COVID-19 pandemic (Bracher et al., 2021); equivalent scores had also been used in previous forecast evaluation efforts (e.g., Hong et al., 2016). The WIS is a proper scoring rule for forecasts that use a set of quantiles to represent a probabilistic forecast distribution. Scores are calculated as a weighted sum of interval scores at different probability levels (e.g., 50% prediction intervals, 80% PIs, 95% PIs, etc…). Larger interval scores indicate less skillful forecasts. An interval score consists of (a) the width of the interval, with larger intervals receiving higher scores (higher scores indicate less accuracy), and (b) a penalty if the interval does not cover the eventual observation, which increases the further away the interval is from the observed value. Equivalently, the WIS can also be characterized as a weighted sum of quantile scores for each individual predictive quantile. The quantile score for a particular quantile level assigns an asymmetric penalty to predictions that are too high or too low, with the relative sizes of the penalties set so that in expectation the score is minimized by the given quantile of the distribution. The most commonly used version of WIS is one that uses an equal weighting of all quantile levels, in which case WIS approximates the continuous ranked probability score (CRPS), a commonly used score for probabilistic forecasts. It is important to note that this weighting was proposed because the resulting score approximates the CRPS, and not because it aligned with any particular public health decision-making rationale. For more mathematical detail on the WIS, we point readers to Bracher et al. (2021).
@@ 3.2 Evaluation metrics
As described above, predictions were submitted to the Forecast Hub in the form of a set of 23 quantiles of the predictive distribution. The WIS can be directly calculated from these quantiles. However, our numerical method for calculating the AS (outlined in appendix C) requires a full cumulative distribution function (CDF) for the forecast in each location. For the purpose of this analysis, we imputed CDFs based on the provided quantiles following a procedure detailed in appendix D. Briefly, this involves interpolating the provided quantiles between the lowest (.01) and highest (.99) probability levels using monotonic cubic splines, and then extending outside these levels with normal distribution tails parametrized to match the two lowest and two highest quantiles. We show in appendix E that if this evaluation procedure had been specified prospectively, the resulting score would be proper, but that a post hoc application of this procedure is improper. We use the procedure here to illustrate the properties of the AS, and note that a collaborative forecasting hub interested in using the AS for evaluation could circumvent propriety issues by communicating the definition of the AS to forecasters and then collecting allocations at specified resource levels as part of forecast submissions.
@@ 3.5.3 Metrics were not consistently correlated over time
We can gain some insight into the discordance between MAS and taMWIS by examining the correlation between time-specific AS and MWIS values of the various models in Figure 5. We find, for example, no clear association between the consistent and high MWIS rank of the JHUAPL-SLPHospEns model and its highly variable AS rank. This contrasts with a clearly positive correlation between MWIS and AS ranks in other models such as Karlen-pypm and USC-SI_kJalpha. We also can see in more detail the consistently high performance of the COVIDhub-ensemble on both scores. And while it did not submit forecasts for enough weeks to be included in Table 2, CMU-TimeSeries is an interesting example of a model that performed consistently well on AS but had only middling MWIS ranks.
@@ 3.5.4 Integrated allocation score across values of K
The integrated allocation score (IAS) summarizes allocation scores (AS) across a range of possible values of the constraint ($K$), allowing one to assign higher weights to more likely values of $K$ (Section 2.2.3). IAS was computed for two weight distributions on $K$, one uniform across the entire range and the other a truncated normal distribution centered at 15,000 (the orange and blue shaded areas in Figure 6A, respectively). Both versions of the IAS were correlated with the original AS for $K=15,000$, with the higher correlation coming from the weighting centered at $K=15,000$ (Figure 6B). Model rankings based on the AS and the centered IAS were roughly similar, with the top and bottom three approaches being the same for both scores (Table 1).
@@ 4 Discussion
Probabilistic forecasts of infectious disease outbreaks have typically been evaluated using well-known proper scoring rules such as the LS, CRPS, and WIS. Often models are ranked according to such a score, but without reference to any underlying decision problem from which the score might be derivable as a Bayes scoring rule or be otherwise motivated. We argue that the value of collaborative forecasting efforts, such as the COVID-19 Forecasting Hub, could be enhanced by including evaluations of forecasts with scoring rules that directly consider the success of forecasts in supporting specific public health decisions. As evidence for this possibility, we have demonstrated how tying forecast evaluation to a specific decision problem (e.g., via the allocation score) can yield model rankings that differ substantively from those based a current standard measure of forecast accuracy, the WIS, which does not refer to any decision problem that has been clearly connected to outbreak response. This result aligns with findings in other fields, especially those where the monetary consequences of using a forecast can be clearly juxtaposed with standard accuracy measures of the forecast (Leitch and Tanner, 1991; Murphy, 1993; Cenesizoglu and Timmermann, 2012). Additionally, we show that an existing ensemble forecast approach was the only method to outperform (in terms of this new evaluation method) two benchmark allocation approaches in a hypothetical application, suggesting that there is room for innovation of current epidemiological modeling techniques that might be spurred on by a reorientation of the field toward practical decision problems. For example, in order to surpass benchmark allocation scores, forecasters need to devote effort to capturing the relative magnitude of future resource need across different locations. Success at this task is not, however, directly rewarded by the WIS.
@@
In practice, epidemiological forecasts are often directed at a diverse set of end-users. It may be easy for some of these users to summarize the consequences of how they use a forecast with a numerical loss function, but for others, such as officials deciding how best to update public “situational awareness” of an outbreak, this may be essentially impossible. Even those forecast uses that are easily framed as expected loss minimization may differ enough that no single score would be appropriate for all users. Ideally, targeted forecasting tools could be developed through close collaboration between modelers and public health officials. However, this may only be possible in settings with sufficient staffing on both an analytics and a public health team. Increasingly, collaborative modeling hubs are being used to generate “one-size-fits-all” forecasts for many locations at once. But it could still be beneficial in these settings to expand the set of scoring methods used by the hub to include scores, such as the AS, which are based on specific public health decision problems. This could help end users better understand the value of available forecasts as inputs to their particular decision making contexts. (We note that issues of propriety are raised when forecasts are evaluated with scoring rules for which they were not elicited or via parameters imputed by a hub after submission, both of which are done in our application example. We consider these issues briefly in appendices D and E.)
@@
There are several important limitations to the current work. The allocation score we developed here does not directly account for important considerations such as fairness or equity of allocations, or more broadly, individual and group preference relations that are difficult or even impossible to encode into a loss or utility function. Ambiguity aversion (in the sense of Ellsberg (1961)), for example, seems especially relevant to the use of forecasts in outbreak response, where there can be immense social and political pressure on public health officials to maintain a transparent base of evidence for their policy choices and recommendations. The proposed framework also does not attempt to capture the broader context of decision making. For example, in practice it may be possible to increase the resource constraint $K$ by shifting funding from other disease mitigation measures. Finally, we emphasize that we opted at this stage of development to explicitly rule out situations where a successful epidemiological forecast could lead to policy decisions that change the distribution of the predicted outcome $Y$. Our framework would need considerable enhancements before being applicable to forecast evaluation beyond horizons for which causal feedback can be neglected.
@@
An opportunity for further investigation is to more carefully evaluate whether forecasts add value to standard procedures already in place for making public health decisions. In the context of allocations, such a procedure might extrapolate need from current observed need and population levels (similarly to the two benchmark approaches presented above), with adjustments based on other political or real-world considerations. For example, in many settings public health stakeholders will make decisions after synthesizing information from a variety of quantitative and qualitative sources coupled with expert judgment. The allocation score presented in this work does not directly measure whether a given forecast adds useful information to such an existing decision-making process. While the scoring procedures as presented do not directly address this question, they could be modified (say, by comparison to a baseline model or expert-elicited allocations in the absence of forecast data) to quantify the benefit of using a forecast to inform a specific decision.
@@ Acknowledgements
We address some technical and methodological points from the main text. We begin in section A by defining proper scoring rules and showing that the allocation score is proper. Section B gives a justification for the result that the Bayes act for the allocation problem is given by a vector of quantiles of the forecast distributions in each location at a shared probability level, which was stated in section 2.2.2. We describe the algorithm that we use to compute allocations given a forecast distribution in each location in section C. In section D, we describe the methods for approximating a full distribution from a set of quantiles, implemented in the R package distfromq, that we used to support computation of allocation scores from quantile forecasts that were submitted to the US COVID-19 Forecast Hub for the application in section 3. Finally, in section E we examine implications for propriety of an analysis that uses summaries of forecasts (such as predictive quantiles) rather than the full forecast distributions for computation of allocation scores.
Appendix A Proper scoring rules
@@
A.1 The allocation score is proper